What mass of Z would crystalize out of a solution containing 100g of water that is saturated at 76 degrees celcius, if the solution is cooled to 10 degrees celcius. Upload working


1
Expert's answer
2021-10-07T03:48:23-0400

Let the amount of z present in the 100g solution be x g.

The amount of water present in the saturated solution is (100-x)

solubility of z "=x/(100-x)\times 100"

solubility(100-x)=100x

(100)solubility-x(solubility)=100x

100x+x(solubility)=100(solubility)

"x=(100\times solubility)/(100+solubility)g" at "76^oC"

Note this value will be lower at "10^oC"